# A search for self-lensing binaries with TESS and constraints on their occurrence rate

### v1

@@ 1 Introduction
Several works have since attempted the search (Sorabella et al., 2023, 2024), but to date, no new SLB candidate from TESS has been confirmed. In this work, we describe our search for non-repeating self-lensing pulses in TESS light curves. We also revisit the Kepler-detected SLBs and constrain the occurrence rate of AU-scale WD+MS binaries. In Section 2, we describe properties of our initial TESS sample and in Section 3, our method to identify candidate self-lensing signals. In Section 4, we present one of the most promising candidates from our search. In Section 5, we describe injection and recovery tests to measure the detection efficiency of our algorithm. Using the results of these tests, in Section 6, we calculate occurrence rates of WD companions to solar-type stars in AU-scale orbits. We also compare the occurrence rate to values inferred from the Gaia astrometric sample of WD + MS binaries. Finally, we conclude in Section 7. Spectroscopic follow-up of our best candidates will be described in future work.
@@ 2.2 Dependence of self-lensing signal on stellar and orbital parameters
We do not explicitly exclude evolved stars in our light curve search. However, these stars have large radii and therefore would not produce self-lensing signals that are strong enough to be detected if the compact object is a WD. In Figure 3, we show how the shape of the pulse depends on the lensed star’s mass and orbital period. Here, we model the pulse as having two components (Foreman-Mackey et al. 2016; code available on github [^1] https://github.com/dfm/transit): (1) an eclipse of the luminous star by the WD (dimming), and (2) an inverted transit / self-lensing pulse where the radius of the eclipsing star is set to be the Einstein radius of the white dwarf (brightening). The second component is a valid approximation as long as the Einstein radius is much smaller than the background star (Agol, 2003), which is the case for WD + MS binaries that we are interested in. Here, we assume a WD of mass $0.6\,M_{\odot}$ and radius $0.01\,R_{\odot}$, and consider an edge-on orbit. For the lensed star, we assume that it is on the main-sequence and use the Eker et al. (2018) mass-radius relation. We take constant quadratic limb-darkening coefficients, $u_{1}=0.4$ and $u_{2}=0.3$ (these have little effect on the amplitude and duration). In the left panel of Figure 3, we plot the model light curves for different orbital periods at a fixed MS star mass. In the center and right panels, we plot the amplitude and duration against orbital period for a few masses. We see that the amplitude is higher at longer orbital periods; hence, we do not expect to detect SLBs with multiple pulses within a single TESS sector. It also falls rapidly with increasing MS star mass, corresponding to larger radii. Meanwhile, the duration increases with increasing MS star mass and orbital period.
@@ 3 Light curve search
Figure 7: Examples of common false positives in our sample: (a) Pulses that are not well fit by any model – here, we show five different models that have been fit to a pulse, none of which are a good representation of the data but where a self-lensing model had the highest BIC, (b) Artifacts occuring at the edge of data gaps, (c) Repeating signals with unclear origins, but with periods too short to be self-lensing.
@@ 3.1 Quality cuts
To cut down on the number of candidates further, we selected only pulses with best-fit durations greater than 0.20.20.2\,0.2d, and amplitudes smaller than 0.04 that occurred in light curves of sources with absolute extinction-corrected Gaia G magnitudes $M_{G},0>0.4$. These are conservative cuts based on models of expected self-lensing pulses over a range of WD masses, companion star radii, and orbital periods. For example, much stronger signals (i.e. larger amplitudes) would require lens masses significantly above the $1.4\,M_{\odot}$ limit for a WD or long orbital periods of $\gtrsim 1000\,$d, where the transit probability is low. The lower limit on the G magnitude removes most giants and MS stars with mass $\gtrsim 3M_{\odot}$, where self-lensing pulses are weak (Section 2.2) and which we find present more false positives (see discussion of heartbeat stars below). By removing events with the shortest durations, we filter out most asteroid crossing events, which were found to be significant contaminants by Sorabella et al. (2023) but which – based on our inspection of the full frame images (Brasseur et al., 2019) of our candidates – appear not to contaminate our final sample. We also removed all sources with a total RV variability (rv_amplitude_robust) measured by Gaia to be $<5\,\mbox{km s}^{-1}$ (see discussion below regarding Figure 8; sources with no measured value was kept). While the five Kepler SLBs are too faint to have measured rv_amplitude_robust, they would have satisfied all these cuts if they were brighter. After this step, about $10$ objects remained per sector, leaving a total of $\sim 800$ candidate signals across all sectors.
@@ 3.2 Visual inspection & false-positives
We then inspected every remaining candidate by eye. The most common false positives appear to be heartbeat stars, with repeating pulses on timescales of days to weeks (e.g. Welsh et al., 2011; Fuller, 2017). While we did not follow up such systems and thus cannot confirm their classification, there are several with radial velocity orbital solutions from Gaia DR3 (Gaia Collaboration et al., 2023) where the times of pulses aligned with the expected periastron times, strongly indicating that these are in fact heartbeat stars. These also typically show a dip in the flux before and after the peak. Figure 6 shows an example of one such object. (For more on heartbeat systems from TESS, see recent work by Solanki et al. 2024.) Examples of other common contaminants are shown in Figure 7. These are signals badly fit by all of the models (panel a), increasing flux at the edges of data gaps due to poor de-trending (b), and repeating signals at periods too short for self-lensing and which have shapes distinct from typical heartbeat stars (c).
@@ 3.3 Gold sample
In Figure 8, we plot the maximum RV variability amplitude (i.e. two times the semi-amplitude) and predicted RUWE as a function of orbital period for a $1.0\,M_{\odot}$ star hosting a $0.6\,M_{\odot}$ companion. We assume an edge-on orbit and zero eccentricity. We predicted RUWE using the gaiamock package, which generates mock Gaia epoch astrometry of astrometric binaries following the Gaia scanning law and fits the data with a single-star model (El-Badry et al. 2024, in prep; code available on GitHub [^3] https://github.com/kareemelbadry/gaiamock). We generate 100 realizations of the same binary with the specified parameters at random coordinates and orientations. We shade the region $1\sigma$ away from the median of the distribution at each period and distance. For periods below 1000 days, we expect RV amplitudes $\gtrsim 20\,$km/s and above 50 days, RUWE $\gtrsim 1.4$. Note that we did not place a stricter cut on rv_amplitude_robust as the exact value of this depends on the fraction of an orbit that is observed (as well as the exact orbital and stellar parameters). The results of our RV follow-up will be presented in future work.
In Table 1 of the appendix, we summarize properties of the most promising targets after visual inspection. We include qualitative notes of their key properties, which determine their order of priority for follow-up. We also plot the detected signals and best-fit model of the 12 most promising candidates in Figure 16.
@@ 3.4 Focused search in the Gaia NSS catalog
We also performed a search isolating objects in the Gaia DR3 non-single stars catalog. In many cases, a spectroscopic or astrometric orbital solution can immediately rule out candidates whose mass functions or inclinations are inconsistent with being SLBs. These solutions also provide constraints on the periastron time, period, and eccentricity so it is possible to check whether a candidate signal occurs during the expected time of eclipse (Section 4). Furthermore, fewer observations are needed to constrain the radial velocity curves, and thus orbits, of these systems.
@@
Here, we removed the requirement of a best-fit transit model as well as any cuts based on the amplitude and duration of the pulse. Instead, we allowed periods between 50 and 1000 days, and goodness_of_fit $<4$ (large values indicate a poor fit to the astrometric model). In the case of astrometric binaries, we only kept objects with inclinations within 5 of edge-on (otherwise, we would not expect an eclipse). Once again, we excluded systems with SB2 solutions.
@@
In Table 2, we summarize properties of the best candidates from this search. As many astrometric binaries are ruled out based on their inclinations, all of the candidates listed here have SB1 solutions. While these do not have constraints on the inclination, we can calculate a lower limit on the companion mass which occurs for an edge-on orbit. In the case that the system is indeed a SLB, this lower limit must be close to the true value so we prioritize systems where it is reasonable for a WD companion. The orbital periods vary from $\sim 50$ to $1000$ days which is in the expected range for self-lensing to be detectable. Several systems are in eccentric orbits, with $e\gtrsim 0.3$. While this does not rule them out as SLBs, they are more likely to be heartbeat systems and possibly in tension with being post-interaction WD + MS binaries which have been tidally circularized (though a few such eccentric binaries are present in the Gaia sample; Shahaf et al. 2024; Yamaguchi et al. 2024b).
4 A promising candidate
Here, we briefly describe one of the most promising targets from our search, TIC 55276301. The light curve of this object, and a zoom-in on the detected pulse, is shown on the top and bottom left panel of Figure 9. This object has an SB1 solution in Gaia DR3 (source ID 631849788720040448) with a $358\pm 2$ day orbital period. It was identified in both the main and NSS-focused searches.
By taking the Gaia orbital parameters, we can fit the pulse and get preliminary constraints on the luminous star mass, $M_{\star}$, and inclination, $i$, in the case it is a self-lensing binary. The radii of the stars are estimated using mass-radius relations (Eker et al., 2018; Nauenberg, 1972), the limb-darkening coefficients are fixed ($u_{1}=0.4,u_{2}=0.3$), and the WD mass is from the mass function of the Gaia SB1 orbit along with the two fitted parameters. The resulting best-fit model is plotted in Figure 9, where $M_{\star}=1.08M_{\odot}$ and $i=89.73^{\circ}$ for which $M_{\rm WD}=0.69M_{\odot}$. This value of $M_{\star}$ is reasonable given the absolute magnitude of the star ($M_{G}=3.8$; Section 2.1) and $M_{\rm WD}$ is within the expected range for a WD. In the future, we will carry out a more thorough analysis and provide a detailed description of the fitting procedure. Here, we only aim to communicate that it is possible to have a self-lensing binary with the Gaia orbit which produces the observed pulse.
Radial velocities from our spectroscopic follow-up will allow us to improve constraints on the orbital parameters. In particular, due to its almost circular orbit ($e\sim 0.02$), the periastron time is very poorly determined. On the bottom right panel of Figure 9, we show the dispersion in the expected transit time (i.e. time of conjunction) calculated from the Gaia solution. We see that the central time of the observed pulse is consistent with the Gaia constraints, but the uncertainty is large because the Gaia solution is based on data obtained 6-9 orbits prior to the TESS observation. A stronger constraint will allow us to confirm or rule out if they do in fact agree with each other.
@@ 5.2 Simulating the WD + MS binary population
We also carried out the same injection and recovery test for Kepler long-cadence light curves from quarter 5. A figure showing the results of this test, similar to Figure 13, can be found in Figure 17 in the appendix. We find similar trends with respect to the different stellar and orbital parameters, but about an order of magnitude higher overall detection rate of $\sim 0.25\%$ compared to TESS. This is primarily due to the 4 year baseline of the observations – since we are considering periods of up to 1000 days, all self-lensing binaries will have at least one observed pulse in this time.
@@ 6 Occurrence rates of AU-scale WD + MS binaries
Using the results of our injection and recovery test (Section 5.2), we can estimate the fraction of all stars in AU-scale WD + MS binaries, $f_{\rm AU}$, based on the number of SLBs discovered from our search. The details of the following derivation are described in Green et al. (in prep), but we summarize the key steps here.
@@
$N_{\rm SLB,found}=f_{\rm AU}\sum_{j=1}^{N_{\rm input}}p(\mbox{detect}|\textbf{{Y}})=f_{\rm AU}N_{\rm input}\bar{\mathcal{S}}$ (4)
@@ 7 Conclusion
  • While our sample of TESS-SPOC light curves is made up of a greater number of stars than Kepler observed in total, due to TESS being a smaller telescope with each star having a shorter baseline, we are less likely to observe a self-lensing event (Figure 5). On the other hand, our stars are on average closer and brighter, making them easier to follow up (Figure 2). In the future, looking at all TESS light curves, as opposed to just those from TESS-SPOC, will significantly increase discovery potential.
@@
  • We present one promising candidate from our search, TIC 55276301, which has an SB1 solution from Gaia (Figure 9). We also provide a list of candidates after manual inspection with large RV amplitudes $>10\,\mbox{kms}^{-1}$ and RUWEs $>1.0$, possibly indicative of binarity at the expected orbital periods (Tables 1, 2).
@@ 8 Acknowledgements
TIC source ID note RUWE RV amplitude $M_{G}$ $M_{\star}$ [$M_{\odot}$] NSS $P_{\rm orb}$ [$M_{\odot}$] $e$ $M_{\rm WD}$ [$M_{\odot}$]
55276301 631849788720040448 good fit, reasonable orbital parameters 3.83 29.44 3.81 1.27 SB1 358 0.02 0.75
82324681 657510191090704640 good fit, large RUWE and RV amp 6.04 24.77 4.12 1.19
455471849 5908076786692427392 good fit, large RUWE and RV amp 2.56 17.92 1.97 1.84
394986259 2173303415148175744 good fit, large RUWE and RV amp 2.14 25.67 3.14 1.47
35015820 112075492260192768 good fit, large RUWE and RV amp 1.75 30.06 1.91 1.88
389661615 6563842117244628736 good fit, large RUWE and RV amp 1.43 19.39 4.06 1.21
320193438 669719821039909888 good fit, large RUWE and RV amp 4.39 23.60 3.87 1.26
62897795 674473319043704320 good fit, large RUWE and RV amp 3.81 21.46 3.89 1.25
123256703 5507496456985455488 good fit, RUWE   1, eccentric 1.11 13.13 0.54 3.16 SB1 347 0.36 0.45
272441807 5358624369305080960 good fit, large RV amp 1.01 24.88 1.76 1.97
437884297 3397923420151903104 second smaller peak may be present 3.34 16.06 3.61 1.33
68809563 3408601151229197440 good fit, but low WD mass 2.22 17.70 2.10 1.77 SB1 210 0.13 0.37
63119966 2125813235891371392 good fit, but low WD mass 4.47 11.63 1.75 1.98 SB1 803 0.09 0.34
427680164 2716570701665963776 good fit, but low WD mass 4.94 10.28 2.14 1.75 SB1 542 0.07 0.29
268224676 152589712603374080 shape narrow near peak 5.53 14.17 3.09 1.49
153110800 5597746020670754304 good fit, large RUWE and RV amp 1.23 37.09 1.50 2.18
440752741 3167264263105849856 good fit, slightly enhanced RUWE 1.19 11.63 2.05 1.79
301499451 3814818921862284032 slight dip in flux before peak 9.07 13.79 5.95 0.83
456934065 95626832665758464 slight dip in flux after peak 1.18 12.08 2.17 1.73
91841940 883988283374275328 slight dip in flux before and after 1.18 19.67 1.85 1.92
270024241 3852384252020976256 dip in flux before peak 1.21 11.73 3.25 1.44
425171701 3162720222066331520 fit okay, other variability present 1.32 19.80 1.14 2.50
302187179 679043507844279680 fit okay, noisy, large RUWE and RV amp 1.28 50.28 1.92 1.87
457200146 3232916033679156224 fit okay, large RV amp 1.13 23.49 1.33 2.32
270268300 2641692752928242432 fit okay, large RV amp 1.10 34.76 3.57 1.34
95237001 868699711628440064 fit okay, large RUWE and RV amp 1.80 17.78 1.94 1.86
314487191 4532321418003888128 fit okay, other variability present 1.43 14.47 0.58 3.11
155813746 5692765712460857472 other variability present, large RV amp 1.06 29.87 1.42 2.24
58641598 152266078227690752 possibly double peaked, large RV amp 1.01 44.77 1.22 2.43
468473101 1965726168256488704 pulse a little weak 4.63 21.55 3.40 1.39
229510522 2247655071463769600 pulse a little weak 2.02 17.48 2.08 1.78
394019762 4543783242611185280 pulse a little weak 3.87 37.11 1.39 2.27
350328162 3867690141074708736 pulse a little weak 3.75 17.67 3.98 1.23
247719397 3346301795905002368 pulse a little weak 1.04 17.62 2.38 1.65
65191369 5611129516718902272 pulse a little weak 2.04 47.09 3.35 1.41
@@
61845524 3109952734908875520 pulse a little weak, highly eccentric 2.30 35.12 2.32 1.67 SB1 147 0.43 0.58
182228119 5539762003663045248 lots of variability, peak unclear 2.72 11.81 0.69 2.98
66716968 3636646017518739968 noisy, peak unclear 1.70 30.32 4.38 1.13
247480759 3347433674406262656 fit okay, noisy 1.41 17.70 5.39 0.92
357181267 3888020828331276928 fit okay 1.17 10.18 2.28 1.68
6994030 656028393013698816 fit okay 1.12 22.47 1.40 2.26
230114166 1648830080850407552 fit okay, very large RV amp 1.12 88.01 2.31 1.67
299891691 6688849614297149824 fit not great, step-like 1.49 10.29 2.92 1.52
124508784 3100778405583395072 fit not great, variability 1.48 20.23 1.91 1.87
372526211 3359159141284450560 fit not great 1.07 12.24 2.01 1.82
9293371 3181637010382678400 fit not great 1.00 20.93 1.54 2.15
369000430 4037702779416807552 second peak within the same sector 1.54 11.65 0.75 2.91
346335796 1931251084240611584 second peak within the same sector 1.49 21.95 1.90 1.88 SB1 65 0.41 0.34
336883262 3345244031360083456 HB-like, possibly dynamical tides 1.19 14.45 2.02 1.81
29850601 5390373282936805376 HB-like shape 1.15 34.08 1.85 1.91
80371229 887223458900219648 HB-like shape 1.13 16.26 2.70 1.56
307364849 5272990074390867840 HB-like shape 1.12 14.60 1.56 2.13
268077447 152958804912813952 HB-like shape 1.43 14.69 2.22 1.71
249835273 450854502451428352 HB-like shape 1.11 12.71 2.35 1.66
234632282 3124127982407698560 HB-like shape 1.09 27.43 1.30 2.35
95235327 878271437088970496 HB-like shape 1.04 21.22 2.41 1.64
69847960 3419335859346402176 HB-like shape 1.03 25.93 1.72 2.00
165325570 679034196355166592 HB-like shape 1.03 17.96 1.62 2.08
27540766 656514789469871744 HB-like shape 3.64 15.31 2.25 1.70
462506516 873550786993324928 HB-like shape 3.36 31.62 4.82 1.03
5537207 62157045883611136 HB-like shape 3.19 10.94 5.13 0.97
178217555 663915563579359616 HB-like shape 2.87 12.03 4.01 1.22
202954782 338858355423218688 HB-like shape 2.82 30.61 4.09 1.20
241234217 971610564013049472 HB-like shape 48.82 30.49 1.34 2.32
55397180 594063727305795584 HB-like shape 13.55 12.81 3.90 1.25
151937435 5397005399634166784 HB-like shape 4.22 25.52 3.14 1.47
453368802 976007747235767168 near edge of data gap 1.59 12.52 1.78 1.96
293164329 4775452026208341248 near edge of data gap 10.64 15.36 5.58 0.89 AstroSpectroSB1 584 0.28 0.34
238004509 5498660609742310272 near edge of data gap 1.00 16.88 0.42 3.32
385110789 2590034024458994304 inclination is too low 3.25 19.91 2.73 1.56 AstroSpectroSB1 279 0.08 0.62
61143248 44958893021221248 inclination is too low 2.60 12.19 2.86 1.54 AstroSpectroSB1 516 0.26 0.28
320193939 669027128714437120 period too short 1.17 42.39 4.16 1.18 SB1 15 0.02 0.31

### v2

@@ 1 Introduction
Several works have since attempted the search (Sorabella et al., 2023, 2024), but to date, no new SLB candidate from TESS has been confirmed. In this work, we describe our search for non-repeating self-lensing pulses in TESS light curves. We also revisit the Kepler-detected SLBs and constrain the occurrence rate of AU-scale WD+MS binaries. In Section 2, we describe properties of our initial TESS sample and in Section 3, our method to identify candidate self-lensing signals. In Section 4, we describe injection and recovery tests to measure the detection efficiency of our algorithm. Using the results of these tests, in Section 5, we calculate occurrence rates of WD companions to solar-type stars in AU-scale orbits. We also compare the occurrence rate to values inferred from the Gaia astrometric sample of WD + MS binaries. Finally, we conclude in Section 6. Spectroscopic follow-up of our best candidates will be described in future work.
@@ 2.2 Dependence of self-lensing signal on stellar and orbital parameters
We do not explicitly exclude evolved stars in our light curve search. However, these stars have large radii and therefore would not produce self-lensing signals that are strong enough to be detected if the compact object is a WD. In Figure 3, we show how the shape of the pulse depends on the lensed star’s mass and orbital period. Here, we model the pulse as having two components (Foreman-Mackey et al. 2016; code available on github [^1] https://github.com/dfm/transit): (1) an eclipse of the luminous star by the WD (dimming), and (2) an inverted transit / self-lensing pulse where the radius of the eclipsing star is set to be $\sqrt{2}$ times the Einstein radius of the white dwarf (brightening). The second component is a valid approximation as long as the Einstein radius is much smaller than the background star (Agol, 2003), which is the case for WD + MS binaries that we are interested in. Here, we assume a WD of mass $0.6\,M_{\odot}$ and radius $0.01\,R_{\odot}$, and consider an edge-on orbit. For the lensed star, we assume that it is on the main-sequence and use the Eker et al. (2018) mass-radius relation. We take constant quadratic limb-darkening coefficients, $u_{1}=0.4$ and $u_{2}=0.3$ (these have little effect on the amplitude and duration). In the left panel of Figure 3, we plot the model light curves for different orbital periods at a fixed MS star mass. In the center and right panels, we plot the amplitude and duration against orbital period for a few masses. We see that the amplitude is higher at longer orbital periods; hence, we do not expect to detect SLBs with multiple pulses within a single TESS sector. It also falls rapidly with increasing MS star mass, corresponding to larger radii. Meanwhile, the duration increases with increasing MS star mass and orbital period.
@@ 3 Light curve search
Figure 7: Examples of common false positives in our sample (other than heartbeats; Figure 6): (a) Pulses that are not well fit by any model – here, we show five different models that have been fit to a pulse, none of which are a good representation of the data but where a self-lensing model had the highest BIC, (b) Artifacts occurring at the edge of data gaps, (c) Repeating signals with unclear origins, but with periods too short to be self-lensing, (d) Asteroid crossing events, as seen in FFIs taken around the mid-pulse time of the central source. (e) Over-subtraction of background eclipses. The light curve in the middle shows dimming of a nearby star (selected pixels boxed in red in the FFI) to the central source which occurs at the same time as the detected pulse.
@@ 3.1 Quality cuts
To cut down on the number of candidates further, we selected only pulses with best-fit durations greater than 0.20.20.2\,0.2d, and amplitudes smaller than 0.04 that occurred in light curves of sources with absolute extinction-corrected Gaia G magnitudes $M_{G},0>0.4$. These are conservative cuts based on models of expected self-lensing pulses over a range of WD masses, companion star radii, and orbital periods. For example, much stronger signals (i.e. larger amplitudes) would require lens masses significantly above the $1.4\,M_{\odot}$ limit for a WD or long orbital periods of $\gtrsim 1000\,$d, where the transit probability is low. The lower limit on the G magnitude removes most giants and MS stars with mass $\gtrsim 3M_{\odot}$, where self-lensing pulses are weak (Section 2.2) and which we find present more false positives (see discussion of heartbeat stars below). We also removed all sources with a total RV variability (rv_amplitude_robust) measured by Gaia to be $<5\,\mbox{km s}^{-1}$ (see discussion below regarding Figure 8; sources with no measured value was kept). While the five Kepler SLBs are too faint to have measured rv_amplitude_robust, they would have satisfied all these cuts if they were brighter. After this step, about $10$ objects remained per sector, leaving a total of $\sim 800$ candidate signals across all sectors.
@@ 3.2 Visual inspection & false-positives
We then inspected every remaining candidate by eye. The most common false positives appear to be heartbeat stars, with repeating pulses on timescales of days to weeks (e.g. Welsh et al., 2011; Fuller, 2017). While we did not follow up such systems and thus cannot confirm their classification, there are several with radial velocity orbital solutions from Gaia DR3 (Gaia Collaboration et al., 2023) where the times of pulses aligned with the expected periastron times, strongly indicating that these are in fact heartbeat stars. These also typically show a dip in the flux before and after the peak. Figure 6 shows an example of one such object. (For more on heartbeat systems from TESS, see recent work by Solanki et al. 2024.) Examples of other common contaminants are shown in Figure 7. These are signals badly fit by all of the models (panel a), increasing flux at the edges of data gaps due to poor de-trending (b), repeating signals at periods too short for self-lensing and which have shapes distinct from typical heartbeat stars (c), and asteroid crossing events (also found to be significant contaminants by Sorabella et al. (2023); a detailed description of one such false positive can be found in Appendix A) (d). To identify and remove possible asteroid crossing events, we inspected FFIs (Brasseur et al., 2019) $\sim 1$ day around the central time of the peak to look for faint sources moving across the camera and traveling over the source. FFIs also revealed several cases of likely background eclipsing binaries, seen as a dimming in a source located close to the target, which can produce a positive pulse in surrounding pixels due to background over-subtraction (e).
@@ 3.3 Gold sample
In Figure 8, we plot the maximum RV variability amplitude (i.e. two times the semi-amplitude) and predicted RUWE as a function of orbital period for a $1.0\,M_{\odot}$ star hosting a $0.6\,M_{\odot}$ companion. We assume an edge-on orbit and zero eccentricity. We predicted RUWE using the gaiamock package, which generates mock Gaia epoch astrometry of astrometric binaries following the Gaia scanning law and fits the data with a single-star model (El-Badry et al. 2024; code available on GitHub [^3] https://github.com/kareemelbadry/gaiamock). We generate 100 realizations of the same binary with the specified parameters at random coordinates and orientations. We shade the region $1\sigma$ away from the median of the distribution at each period and distance. For periods below 1000 days, we expect RV amplitudes $\gtrsim 20\,$km/s and above 50 days, RUWE $\gtrsim 1.4$. Note that we did not place a stricter cut on rv_amplitude_robust as the exact value of this depends on the fraction of an orbit that is observed (as well as the exact orbital and stellar parameters). The results of our RV follow-up will be presented in future work.
In Table 1 of Appendix B, we summarize properties of the most promising targets after visual inspection. We include qualitative notes of their key properties, which determine their order of priority for follow-up. We also plot the detected signals and best-fit model of the 12 most promising candidates in Figure 16.
@@ 3.4 Focused search in the Gaia NSS catalog
We also performed a search isolating objects in the Gaia DR3 non-single stars catalog. In many cases, a spectroscopic or astrometric orbital solution can immediately rule out candidates whose mass functions or inclinations are inconsistent with being SLBs. These solutions also provide constraints on the periastron time, period, and eccentricity so it is possible to check whether a candidate signal occurs during the expected time of eclipse. Furthermore, fewer observations are needed to constrain the radial velocity curves, and thus orbits, of these systems.
@@
Here, we removed the requirement of a best-fit transit model as well as any cuts based on the amplitude and duration of the pulse. Instead, we allowed periods between 50 and 1000 days which is in the expected range for self-lensing to be detectable, and goodness_of_fit $<4$ (large values indicate a poor fit to the astrometric model). In the case of astrometric binaries, we only kept objects with inclinations within 5 of edge-on (otherwise, we would not expect an eclipse). Once again, we excluded systems with SB2 solutions.
@@
Many astrometric binaries were ruled out based on their inclinations, and others were identified to be asteroid crossing events from their FFIs. We include one additional candidate from this search in Table 1. A few of our candidates have SB1 solutions. While SB1s do not have constraints on the inclination, we can calculate a lower limit on the companion mass which occurs for an edge-on orbit. In the case that the system is indeed a SLB, this lower limit must be close to the true value. We removed several systems in eccentric orbits, with $e\gtrsim 0.3$. While this does not rule them out as SLBs, they are more likely to be heartbeat systems and possibly in tension with being post-interaction WD + MS binaries which have been tidally circularized (though a few such eccentric binaries are present in the Gaia sample; Shahaf et al. 2024; Yamaguchi et al. 2024b).
@@ 4.2 Simulating the WD + MS binary population
We also carried out the same injection and recovery test for Kepler long-cadence light curves from quarter 5. A figure showing the results of this test, similar to Figure 12, can be found in Figure 17 in Appendix C. We find similar trends with respect to the different stellar and orbital parameters, but about an order of magnitude higher overall detection rate of $\sim 0.25\%$ compared to TESS. This is primarily due to the 4 year baseline of the observations – since we are considering periods of up to 1000 days, all self-lensing binaries will have at least one observed pulse in this time.
@@ 5 Occurrence rates of AU-scale WD + MS binaries
Using the results of our injection and recovery test (Section 4.2), we can estimate the fraction of all stars in AU-scale WD + MS binaries, $f_{\rm AU}$, based on the number of SLBs discovered from our search. The details of the following derivation are described in Green et al. (2024), but we summarize the key steps here.
@@
[…]
@@ 6 Conclusion
  • While our sample of TESS-SPOC light curves is made up of a greater number of stars than Kepler observed in total, due to TESS being a smaller telescope with each star having a shorter baseline, we are less likely to observe a self-lensing event (Figure 5). On the other hand, our stars are on average closer and brighter, making them easier to follow up (Figure 2). In the future, we will search through all TESS light curves, as opposed to just those from TESS-SPOC, which is expected to significantly increase discovery potential.
@@
  • We provide a list of candidates after manual inspection with large RV amplitudes $>10\,\mbox{kms}^{-1}$ and RUWEs $>1.0$, possibly indicative of binarity at the expected orbital periods (Table 1).
@@ 7 Acknowledgements
Appendix A A false positive
In an earlier version of our manuscript, we described a promising target from our search, TIC 55276301. After taking a closer look at its FFIs, it has been identified to be a faint asteroid crossing event. Nevertheless, we leave a description of its properties here for reference.
The light curve of this object, and a zoom-in on the detected pulse, is shown on the top and bottom left panel of Figure 15. This object has an SB1 solution in Gaia DR3 (source ID 631849788720040448) with a $358\pm 2$ day orbital period. It was identified in both the main and NSS-focused searches.
By taking the Gaia orbital parameters, we can fit the pulse and get preliminary constraints on the luminous star mass, $M_{\star}$, and inclination, $i$, in the case it is a self-lensing binary. The radii of the stars are estimated using mass-radius relations (Eker et al., 2018; Nauenberg, 1972), the limb-darkening coefficients are fixed ($u_{1}=0.4,u_{2}=0.3$), and the WD mass is from the mass function of the Gaia SB1 orbit along with the two fitted parameters. The resulting best-fit model is plotted in Figure 15, where $M_{\star}=1.08M_{\odot}$ and $i=89.73^{\circ}$ for which $M_{\rm WD}=0.69M_{\odot}$. This value of $M_{\star}$ is reasonable given the absolute magnitude of the star ($M_{G}=3.8$; Section 2.1) and $M_{\rm WD}$ is within the expected range for a WD.
Due to its almost circular orbit ($e\sim 0.02$), the periastron time is very poorly determined. On the bottom right panel of Figure 15, we show the dispersion in the expected transit time (i.e. time of conjunction) calculated from the Gaia solution. We see that the central time of the observed pulse is consistent with the Gaia constraints, but the uncertainty is large because the Gaia solution is based on data obtained 6-9 orbits prior to the TESS observation.
While we have ruled out this source to be a SLB, we note that follow-up spectra show no contribution from a second luminous companion which would be expected for a MS star of mass $\gtrsim 0.7\,M_{\odot}$. Therefore, it is likely a WD + MS binary, though not one that is self-lensing.
Appendix B List of SLB candidates
In Table 1, we list candidate SLBs from our main and focused NSS searches, and in Figure 16, we show the light curves of a subset of these around the detected peaks (Section 3).
Appendix C Kepler Injection and Recovery
In Figure 17, we show the results of performing injection and recovery tests on Kepler, compared to TESS, light curves (Section 4.2). A detailed description of each panel can be found in the caption of Figure 12.
@@ Appendix C Kepler Injection and Recovery
TIC source ID Search Note RUWE RV amplitude $M_{G}$ $M_{\star}$ [$M_{\odot}$] NSS $P_{\rm orb}$ [$M_{\odot}$] $e$ $M_{\rm WD}$ [$M_{\odot}$]
82324681 657510191090704640 main Good fit, large RUWE and RV amp, dip in Sector 46 caused by asteroid 6.04 24.77 4.12 1.19 nan nan nan nan
320193438 669719821039909888 main Good fit, large RUWE and RV amp, short duration 4.39 23.60 3.87 1.26 nan nan nan nan
@@
247719397 3346301795905002368 main Pulse a little weak, somewhat crowded, highly eccentric 1.04 17.62 2.38 1.65 nan nan nan nan
61845524 3109952734908875520 main Pulse a little weak, somewhat crowded 2.30 35.12 2.32 1.67 SB1 147 0.43 0.58
457200146 3232916033679156224 main Fit okay, large RV amp, but HB-like shape 1.13 23.49 1.33 2.32 nan nan nan nan
314487191 4532321418003888128 main Fit okay, other variabiliy present 1.43 14.47 0.58 3.11 nan nan nan nan
155813746 5692765712460857472 main Other variability present, large RV amp, very crowded 1.06 29.87 1.42 2.24 nan nan nan nan
468473101 1965726168256488704 main Pulse a little weak, very crowded 4.63 21.55 3.40 1.39 nan nan nan nan
229510522 2247655071463769600 main Pulse a little weak, CVZ but only one pulse 2.02 17.48 2.08 1.78 nan nan nan nan
394019762 4543783242611185280 main Pulse a little weak, crowded 3.87 37.11 1.39 2.27 nan nan nan nan
65191369 5611129516718902272 main Pulse a little weak, very crowded 2.04 47.09 3.35 1.41 nan nan nan nan
247480759 3347433674406262656 main Fit okay, noisy, crowded, signal appears off source 1.41 17.70 5.39 0.92 nan nan nan nan
230114166 1648830080850407552 main Fit okay, very large RV amp 1.12 88.01 2.31 1.67 nan nan nan nan
124508784 3100778405583395072 main Fit not great, variability, crowded 1.48 20.23 1.91 1.87 nan nan nan nan
299891691 6688849614297149824 main Fit not great, step-like 1.49 10.29 2.92 1.52 nan nan nan nan
9293371 3181637010382678400 main Fit not great 1.00 20.93 1.54 2.15 nan nan nan nan
369000430 4037702779416807552 main Two peaks but slightly different relative flux 1.54 11.65 0.75 2.91 nan nan nan nan
29850601 5390373282936805376 main HB-like shape 1.15 34.08 1.85 1.91 nan nan nan nan
307364849 5272990074390867840 main HB-like shape 1.12 14.60 1.56 2.13 nan nan nan nan
234632282 3124127982407698560 main HB-like shape 1.09 27.43 1.30 2.35 nan nan nan nan
151937435 5397005399634166784 main HB-like shape, crowded field 4.22 25.52 3.14 1.47 nan nan nan nan
453368802 976007747235767168 main near edge of data gap 1.59 12.52 1.78 1.96 nan nan nan nan
293164329 4775452026208341248 main near edge of data gap 10.64 15.36 5.58 0.89 ASB1 584 0.28 0.34
238004509 5498660609742310272 main near edge of data gap 1.00 16.88 0.42 3.32 nan nan nan nan
320193939 669027128714437120 main short pulse duration, period too short 1.17 42.39 4.16 1.18 SB1 15 0.02 0.31
63119966 2125813235891371392 nss good fit, low WD mass, possibly background eclipse 4.47 11.63 1.75 1.98 SB1 803 0.09 0.34